If sample space S={e1,e2,e3,e4}, find P(e1) if P(e2)=1/5 and P(e3)=1/3
1
Expert's answer
2021-07-14T15:49:43-0400

"S=\{e_1,e_2,e_3,e_4\}"


Given,

"P(e_2)=1/5\\P(e_3)=1/3"


and we know that

"P(e_1)+P(e_2)+P(e_3)+P(e_4)=1\\\ \\\implies P(e_1)+\dfrac{1}{5}+\dfrac{1}{3}+P(e_4)=1\\\ \\\implies P(e_1)+P(e_4)=1-\dfrac{1}{5}-\dfrac{1}{3}=\dfrac{7}{15}"


Let "P(e_4)=\dfrac{1}{15}"


So, "P(e_1)=\dfrac{7}{15}-\dfrac{1}{15}=\dfrac{6}{15}"